OpenStudy (k8lyn911):

Trigonometry question: Solve the following equation for all radian solutions and if 0 ≤ t < 2π. Give all answers as exact values in radians. (Enter your answers as a comma-separated list. If there is no solution, enter NO SOLUTION.) 5 sin t + 8 = −3 sin t *I think I'm doing it wrong. How I tried to solve it: -3 sin t = -8 sin t = 8/3 8/3 isn't in the trig unit circle.

zepdrix (zepdrix):

Hmm I'm confused by your steps, the 5sin t disappeared somehow :o

zepdrix (zepdrix):

So what you want to do is combine like terms (adding 3sin t to each side),\[\Large\rm 8\sin t+8=0\]Then isolate the sin t by doing a few algebra steps,

OpenStudy (k8lyn911):

That makes a lot more sense. (The video I watched to figure it out wasn't explained, just worked out.) 8 sin t + 8 = 0 8 sin t = -8 sin t = -8/8 So sin t = -1 and t = 3pi/2

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (k8lyn911):

Thanks! :)